Peter is flying a kite using a 20 m long strip. His hand and the kite are 1 m and 15 m above the ground, respectively. Find the angle of depression of Peter from the kite.


1
Expert's answer
2021-04-14T12:46:38-0400

Length of strip of kite = 20m

Height from his hand to kite = 14m

Let, angle of depression of Peter from the kite = "\alpha"

Using sin "\alpha" = perpendicular / hypotenuse

here, sin"\alpha = 14/20"

sin"\alpha = 0.7"

"\alpha = \arcsin(0.7)"

"\alpha = 44.4 \degree"

Angle of depression = "44.4 \degree"